OpenStudy (anonymous):

Need help finding f(f(x)) from f(x) = x^2 + 1 I will solve it, I just need help.

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

I know that I have to solve from the inside out, but I don't understand how to find f(x) when it says f(x) = x^2 + 1

OpenStudy (anonymous):

would it be x^4 + 2?

OpenStudy (anonymous):

You are not being asked to solve the equation. You are being asked to create a new equation f(f(x))=etc, knowing that the rule for the f() function is f(whatever) =(whatever)^2 + 1

OpenStudy (mathmale):

I like to re-write functions such as f(x) = x^2 + 1 as f( ) = ( )^2 + 1 when calculating composites. f(f(x)) is simply the function f(x) with f(x) substituted for x: f(f(x)). Then my f( ) = ( )^2 + 1 becomes f(f(x)) = (x^2+1)^2 + 1. Please do the algebra and simplify the result.

OpenStudy (loser66):

That's why I delete my comments, hehehe. others' instructions are always better than mine, Thanks @douglaswinslowcooper and @mathmale